OpenStudy (anonymous):

A jar of dimes and quarters contains $15.25. There are 103 coins in all. How many of each are there?

OpenStudy (anonymous):

how many unknowns do you have ?

OpenStudy (anonymous):

2, number of dimes and number of quarters are unknown.

OpenStudy (skullpatrol):

How many facts do you have?

OpenStudy (anonymous):

2 facts, we know 15.25 is the total value of the coins and we know that there are 103 coins

OpenStudy (skullpatrol):

Choose a unknown and use it with a give fact to represent the other unknown.

OpenStudy (anonymous):

q = 1525 - d

OpenStudy (skullpatrol):

Let d=number of dimes q= number of quarters Choose an unknown, d or q, and use it with a given fact, d + q = 103, to represent the other unknown.

OpenStudy (anonymous):

q = 103 - d ?

OpenStudy (skullpatrol):

Correct.

OpenStudy (skullpatrol):

Reread the problem and write an equation that represents relationships among the numbers in the problem.

OpenStudy (skullpatrol):

Here, we will use the other given fact that the total value of coins is $15.25.

OpenStudy (skullpatrol):

Remember: d=number of dimes q= number of quarters

OpenStudy (anonymous):

25(103 - d) + 10d = 15.25?

OpenStudy (skullpatrol):

Sooo close.. . just write $15.25 as 1525 because 10cents/dime * number of dimes = number of cents right?

OpenStudy (anonymous):

oh alright, so 25(103-d) + 10d = 1525

OpenStudy (skullpatrol):

Solve the equation and find the unknowns asked for.

OpenStudy (skullpatrol):

"How many of each are there?"

OpenStudy (anonymous):

d = 70 q = 33?

OpenStudy (skullpatrol):

Check your results with the words of the problem. Give the answer.

OpenStudy (anonymous):

yep its correct, thank you for the help

OpenStudy (skullpatrol):

np :)